# Point configurations in sets of sufficient topological structure and a topological Erdős similarity conjecture

### v1

@@ Abstract.
We explore the occurrence of point configurations within non-meager (second category) Baire sets. A celebrated result of Steinhaus asserts that $A+B$ and $A-B$ contain an interval whenever $A$ and $B$ are sets of positive Lebesgue measure in $\mathbb{R}^{n}$ for $n\geq 1$. A topological analogue attributed to Piccard asserts that both $AB$ and $AB^{-1}$ contain an interval when $A,B$ are non-meager (second category) Baire sets in a topological group. We explore generalizations of Piccard’s result to more complex point configurations and more abstract spaces. In the Euclidean setting, we show that if $A\subset\mathbb{R}^{d}$ is a non-meager Baire set and $F=\{x_{n}\}_{n\in\mathbb{N}}$ is a bounded sequence, then there is an interval of scalings $t$ for which $tF+z\subset A$ for some $z\in\mathbb{R}^{d}$. That is, the set
$\Delta_{F}(A)=\{t\in\mathbb{R}:\exists z\text{ such that }tF+z\subset A\}$
has nonempty interior. More generally, if $V$ is a topological vector space and $F=\{x_{n}\}_{n\in\mathbb{N}}\subset V$ is a bounded sequence, we show that if $A\subset V$ is non-meager and Baire, then $\Delta_{F}(A)$ has nonempty interior. The notion of boundedness in this context is described below. Note that the sequence $F$ can be countably infinite, which distinguishes this result from its measure-theoretic analogue. In the context of the topological version of Erdős’ similarity conjecture, we show that bounded countable sets are universal in non-meager Baire sets.
@@ 1. Introduction
A large body of research concerns the occurrence (or absence) of patterns in sets of sufficient Hausdorff dimension, see for instance, [5, 6, 7, 15, 19], or in sets of sufficient Fourier dimension [3, 1, 24]. Another notion of size that has proved useful in the study of point configurations is that of Newhouse thickness, see for instance, [16, 17, 27, 23]. In this article, we focus on the occurrence of infinite point configurations within sets of sufficient topological structure.
Our story begins with a well-known theorem of Steinhaus from the 1920s and its topological analogue, known as Piccard’s theorem. Let $A,B\subset\mathbb{R}^{n}$ denote sets of positive measure. Steinhaus’ theorem states that the set $A-B$ contains a nontrivial open set (see [26] for $n=1$ and [12] for $n\geq 1$). Said differently, there is an open set of shifts $t$ so that
@@
This assertion can be proved by observing that the convolution of the characteristic functions $\chi_{A}$ and $\chi_{B}$ is a continuous function, so that the function
@@
For more intricate configurations, these same proof techniques can be recycled to show that for each $k\in\mathbb{N}$, there exists an open ball $S_{k}$ about the origin so that if $\{v_{i}\}_{i=1}^{k}\subset S_{k}$, then
(2) $\bigcap_{i=1}^{k}(A+v_{i})\neq\varnothing.$
@@
It is an immediate consequence that, if $F=\{v_{i}\}_{i=1}^{k}$ denotes an arbitrary finite point configuration, then $A$ contains all sufficiently small affine copies of $F$. That is, if
$\Delta_{F}(A)=\{t\in\mathbb{R}:\exists z\in\mathbb{R}^{n}\ tF+z\subset A\},$
then $\Delta_{F}(A)$ contains an interval with left endpoint 00.
@@
We note that there are many further generalizations and modifications of Steinhaus’ theorem, including a result of Erdős and Oxtoby showing that if $F$ on $\mathbb{R}\times\mathbb{R}$ is a continuously differentiable real-valued function with non-vanishing partial derivatives on its domain, then $F(A,B)$ has nonempty interior [2]. This theorem also holds in higher dimensions, see [8, Chapter II] and the references therein and [25, Proposition 2.9 (i)].
@@
A topological analogue of Steinhaus’ theorem known as Piccard’s theorem asserts that the sum of two non-meager Baire sets (see Definition 1.1) in a topological vector space contains a nonempty open set [21, 13]. Again, addition can be replaced by a two variable function satisfying minimal differentiability assumptions, see [25, Proposition 2.9 (iii)] and [8, Chapter II]. Also see [22] for an analogous statement for topological groups, which states that $A*B$ (and $A*B^{-1}$) contains a nonempty open set provided that $A$ and $B$ are non-meager Baire sets.
While the aforementioned results consider operations on pairs of points, the present article focuses on general point configurations. This work fits into a large and expanding literature on the presence of geometric patterns in fractal sets. The basic problem is to understand how large a subset of Euclidean space must be to ensure that it contains the vertices of a congruent and possibly scaled copy of a given polyhedron or another geometric shape.
This article fits into the broader context of investigating a topological variant of the Erdős similarity conjecture:
For each uncountable set $F$, there exists a non-meager Baire set $A$ that contains no affine image of $F$.
We briefly summarize what is known about this conjecture. Let $E\subset\mathbb{R}^{d}$, and let $X$ denote a collection of subsets in $\mathbb{R}^{d}$. We say that $E$ is universal in $X$ if every $K\in X$ contains an affine copy of $E$.
Universality can be considered for different classes of sets. In the setting of the classic Erdős similarity conjecture, there is a distinction between considering universality over full measure sets versus universality over measurable sets. This distinction persists for the topological Erdős similarity conjecture with co-meager sets and non-meager sets in place of full measure and positive measure respectively. An example of a set that is non-meager but not co-meager is $[0,1]$; more generally, any set with both nonempty interior and exterior is a non-meager set which is not co-meager. For the full story on the Erdős similarity conjecture, see [10].
It is a straightforward consequence of the Baire category theorem that countable sets are universal in the collection of dense $G_{\delta}$ sets (see Example A.9 in Section A). Further, since every co-meager set contains a dense $G_{\delta}$ set, it follows that the countable sets are universal in the collection of co-meager sets.
@@
Our main Euclidean result (Theorem 2.1) extends this observation to non-meager Baire sets; we show that bounded countable sets are universal in non-meager Baire sets. Further, we prove a more quantitative result about an interval worth of scalings, and our results extend to topological vector spaces (Theorem 3.3).
For uncountable sets, Gallagher, Lai, and Weber [4] proved that certain Cantor sets are not universal in non-meager Baire sets, so that Theorem 2.1 need not generically hold with countable replaced by uncountable (see Example A.8 in Section A. Also see, [9] where Jung and Lai demonstrate that a set is universal in the collection of dense $G_{\delta}$ sets if and only if it is a strong measure zero set.
@@ 1.1. Background and definitions
Let $X$ denote a topological space. A subset of $X$ is meager (first category) if it is a countable union of nowhere dense sets. A set is non-meager (second category) if it is not meager. Finally, a set $A\subset X$ is said to be Baire if there is an open set $U$ and a meager set $P$ so that $A$ equals the symmetric difference $A=U\Delta P$.
@@
Equivalently, one may define Baire sets as sets $A$ for which there exists an open set $U$ such that $U\Delta A$ is meager. If $A$ satisfies this condition, then $A$ satisfies Definition 1.1 with $P=U\Delta A$. Conversely, if $U$ satisfies Definition 1.1, then $A\Delta U=P$ is meager. See Theorem A.1 for other useful equivalent definitions.
Morally, one may think of meager sets as the topological analogues of measure zero sets (in the sense that they are considered negligible and are stable under countable unions), and of Baire sets as the topological analogues of measurable sets (in the sense that they are “almost” equal to open sets up to negligible error). A word of caution: even meager sets can have full measure. A succinct account of basic facts and examples is given in Appendix A; for a more thorough introduction, see [11, Chapter I.8].
@@
The definition of $\Delta_{F}(A)$ given above immediately extends to vector spaces:
@@
Let $V$ be a real vector space. Given sets $A,F\subset V$, define
$\Delta_{F}(A)=\{t\in\mathbb{R}:\exists z\in V\ tF+z\subset A\}.$
In the notation above, a consequence of Piccard’s Theorem is that $\Delta_{F}(A)$ has nonempty interior whenever $A$ is a non-meager Baire set and $F$ is a two-point set. In the following sections, we explore generalizations of this result to more complex point configurations and more abstract spaces. To make sense of the set $\Delta_{F}(A)$, we need both translations and scalings, and for this reason we work in topological vector spaces.
@@ 1.2. Acknowledgment
K.T. is supported in part by the Simons Foundation Grant GR137264. Both authors thank Angel Cruz, Yeonwook Jung, and Yuveshen Mooroogen for interesting discussions related to this article.
@@ 2. Results in Euclidean space
Let $A\subset\mathbb{R}^{d}$ be a non-meager Baire set, and let $F=\{x_{n}\}_{n\in\mathbb{N}}$ be a bounded sequence. Then, $\Delta_{F}(A)$ has nonempty interior.
@@
Since $A$ is Baire, there is an open set $U$ such that $A\Delta U$ is meager (since $A=U\Delta P$ implies $A\Delta U\subset P$). Since $A$ is non-meager, $U$ must be nonempty. Let $B$ be any open ball contained in $U$. Since
@@
is meager. If $A\cap B_{t}$ were meager, it would follow that $(I\cap J_{t})$ is meager. This cannot happen if $t$ is chosen so that $I\cap J_{t}$ is a non-degenerate interval. Therefore, for an interval worth of $t$, we have $A\cap B_{t}\neq\varnothing$. Equivalently, $A+B$ contains an interval. ∎
This proof idea generalizes to arbitrary countable configurations without too much difficulty. The key geometric fact we need to use is that the intersection of infinitely many balls still has nonempty interior, provided that each ball is a small translate of a single original ball.
@@
Define $A_{n,t}=A-tx_{n}$. We have $\{tx_{n}+z:n\in\mathbb{N}\}\subset A$ if and only if $z\in\bigcap_{n=1}^{\infty}A_{n,t}$. We must show this intersection is nonempty for an interval worth of $t$. Apply Lemma 2.2 to obtain an open ball $B$ with $B\setminus A$ meager, and let $B_{n,t}=B-tx_{n}$. It follows that $B_{n,t}\setminus A_{n,t}$ is meager for every $n,t$. Finally, let
$\widetilde{B}_{t}=\bigcap_{n=1}^{\infty}B_{n,t}.$
The set $\widetilde{B}_{t}\setminus A_{n,t}$ is meager, since $\widetilde{B}_{t}\subset B_{n,t}$. Therefore,
$\widetilde{B}_{t}\setminus\left(\bigcap_{n=1}^{\infty}A_{n,t}\right)=\bigcup_{n=1}^{\infty}(\widetilde{B}_{t}\setminus A_{n,t})$
is meager. If $\bigcap_{n=1}^{\infty}A_{n,t}$ were meager, then $\widetilde{B}_{t}$ would be meager. However, if $t$ is sufficiently small (relative to the radius of $B$ and $\sup_{n}|x_{n}|$), then each set $B_{n,t}$ is a small translate of the original ball $B$, and hence the intersection $\widetilde{B}_{t}$ has nonempty interior. Therefore, the theorem holds when $J$ is a sufficiently small neighborhood of 00. ∎
@@ 3. Results in topological vector spaces
We establish the claim for $A-B$. Let $U_{1}$, $U_{2}$ denote open sets, and let $P_{1}$, $P_{2}$ denote meager sets so that $A=U_{1}\Delta P_{1}$ and $B=U_{2}\Delta P_{2}$. Let us assume that nonempty open sets are not meager, for otherwise $X$ would also be meager by the Banach category theorem (see Theorem 3.4) as a union of translates of such a set.
@@
(4) $A\cap\left(t+B\right)\supset(U_{1}\cap(t+U_{2}))\backslash(P_{1}\cup(P_{2}+t)),$
and since $(U_{1}\cap(t+U_{2}))$ is a nonempty open set, while $(P_{1}\cup(P_{2}+t))$ is meager, it follows that the right-hand-side of (4) is nonempty.
Moving on, recall that in Theorem 2.1, our hypotheses included that the configuration under consideration was bounded. Since an arbitrary topological vector space need not be metric, we need a definition of boundedness in this context.
@@
Let $V$ be a topological vector space, and let $F\subset V$. We say $F$ is bounded if, for any open set $U\subset V$, there exist $t>0,z\in V$ such that $tF+z\subset U$.
@@
Let $V$ be a topological vector space, and let $F=\{x_{n}\}_{n}\subset V$ be a bounded sequence. If $A\subset V$ is a non-meager Baire set, then $\Delta_{F}(A)$ has nonempty interior.
@@
See [20, Chapter 16] for an exposition of this result. As an immediate consequence, we have:
@@
Next, we need a topological vector space analogue of the open balls of Euclidean space. In Definition 3.2, if $F$ is a bounded set and we scale $F$ by $t_{0}$ to be in a given neighborhood of 0, then there is no guarantee that we could then scale by a smaller $t<t_{0}$ and still guarantee this containment. Fortunately, in any topological vector space, there is a topological base consisting of sets which are closed under scaling down. This is captured in the following lemma.
@@
Let $V$ be a topological vector space, let $U$ be an open set, and let $F$ be a bounded set. For each $x\in F$, define $U_{x,t}=U-tx$. Finally, define
$\widetilde{U}_{t}=\bigcap_{x\in F}U_{x,t}.$
@@
is open. Since Cartesian products of open sets are a basis for the product topology, there exist open sets $U_{1},U_{2}\subset V$ such that $(\textbf{0},\textbf{0})\in U_{1}\times U_{2}$ and $U_{1}+U_{2}\subset U$. The set $U^{\prime}=U_{1}\cap U_{2}\cap U$ is therefore an open neighborhood of 0, contained in $U$, with the property that $U^{\prime}+U^{\prime}\subset U$. By Lemma 3.7, we may take a subset $B\subset U^{\prime}$ with these same properties, which is also balanced. Since $F$ is bounded, there exists $t_{0}$ such that $t_{0}F\subset B$. Since $B$ is balanced, we therefore have $tF\subset B$ for any $t\in J:=(0,t_{0})$. We claim that for any such $t$, the set $\widetilde{U}_{t}$ contains $B$ as a subset. Indeed, for any $z\in B,t\in J$, and $x\in F$, we have
@@
Define $A_{n,t}=A-tx_{n}$. We have $\{tx_{n}+z:n\in\mathbb{N}\}\subset A$ if and only if $z\in\bigcap_{n=1}^{\infty}A_{n,t}$. We must show that this intersection is nonempty for an interval worth of $t$. Let $U$ be an open set and let $P$ be a meager set such that $A=U\Delta P$; let $U_{n,t}=U-tx_{n}$ and $P_{n,t}=P-tx_{n}$. Since any translation of a nowhere dense set is still nowhere dense, the set $P_{n,t}$ is meager for every $n,t$. Define
$\widetilde{U}_{t}=\bigcap_{n=1}^{\infty}U_{n,t}\,\,\,\text{ and }\,\,\,\widetilde{P}_{t}=\bigcup_{n=1}^{\infty}P_{n,t}.$
@@
$\bigcap_{n=1}^{\infty}A_{n,t}\supset\widetilde{U}_{t}\backslash\widetilde{P}_{t}.$
The proof is completed by arguing that $\widetilde{U}_{t}$ is non-meager; indeed, since $\widetilde{P}_{t}$ is meager, then it would follow that $\bigcap_{n=1}^{\infty}A_{n,t}$ is nonempty. By Lemma 3.8, there is a non-degenerate interval $J\subset\mathbb{R}$ such that for any $t\in J$, the set $\widetilde{U}_{t}$ has nonempty interior. By Lemma 3.5, since $V$ contains a non-meager set $A$, a set with nonempty interior must be non-meager.

### v2

@@ Abstract.
We explore the occurrence of point configurations in non-meager Baire sets. A celebrated result of Steinhaus asserts that $A+B$ and $A-B$ contain an interval whenever $A$ and $B$ are sets of positive Lebesgue measure in $\mathbb{R}^{d}$ for $d\geq 1$. A topological analogue attributed to Piccard asserts that both $AB$ and $AB^{-1}$ contain an interval when $A,B$ are non-meager Baire sets in a topological group. We explore generalizations of Piccard’s result to more complex point configurations and more abstract spaces. In the Euclidean setting, we show that if $A\subset\mathbb{R}^{d}$ is a non-meager Baire set and $P=\{v^{i}\}_{i\in\mathbb{N}}$ is a bounded sequence, then there is an interval of scalings $t$ for which $tP+z\subset A$ for some $z\in\mathbb{R}^{d}$. That is, the set
$\Delta_{P}(A)=\{t>0:\exists z\text{ such that }tP+z\subset A\}$
has nonempty interior. More generally, if $V$ is a topological vector space and $P=\{v^{i}\}_{i\in\mathbb{N}}\subset V$ is a bounded sequence, we show that if $A\subset V$ is non-meager and Baire, then $\Delta_{P}(A)$ has nonempty interior. The notion of boundedness in this context is described below. Note that the sequence $P$ can be countably infinite, which distinguishes this result from its measure-theoretic analogue. In the context of the topological version of Erdős’ similarity conjecture, we show that bounded countable sets are universal in non-meager Baire sets.
@@ 1. Introduction
A large body of research concerns the occurrence (or absence) of patterns in sets of sufficient Hausdorff dimension, see for instance, [6, 8, 9, 18, 22], or in sets of sufficient Fourier dimension [4, 1, 27]. Another notion of size that has proved useful in the study of patterns is that of Newhouse thickness, see for instance, [19, 20, 31, 26]. In this article, we focus on the occurrence of infinite point configurations within sets of sufficient topological structure.
Our story begins with a well-known theorem of Steinhaus from the 1920s and its topological analogue, known as Piccard’s theorem. Let $A,B\subset\mathbb{R}^{d}$ denote sets of positive measure. Steinhaus’ theorem states that the set $A-B$ contains a nontrivial open set (see [30] for $d=1$ and [15] for $d\geq 1$). Said differently, there is an open set of shifts $t$ so that
@@
This assertion can be proved by observing that the convolution of the characteristic functions $\chi_{A}$ and $\chi_{B}$ is a continuous, so that the function
@@
For more intricate configurations, these same proof techniques can be recycled to show that for each $k\in\mathbb{N}$, there exists an open ball $S_{k}$ about the origin so that if $\{v^{i}\}_{i=1}^{k}\subset S_{k}$, then
(2) $\bigcap_{i=1}^{k}(A+v^{i})\neq\varnothing.$
@@
It is an immediate consequence that, if $P=\{v^{i}\}_{i=1}^{k}$ denotes an arbitrary finite point set, then $A$ contains all sufficiently small affine copies of $P$. That is, if
$\Delta_{P}(A)=\{t>0:\exists z\in\mathbb{R}^{d}\ tP+z\subset A\},$
then $\Delta_{P}(A)$ contains an interval with left endpoint 00.
@@
There are many further generalizations and modifications of Steinhaus’ theorem, including a result of Erdős and Oxtoby showing that if $F$ on $\mathbb{R}\times\mathbb{R}$ is a continuously differentiable real-valued function with non-vanishing partial derivatives on its domain, then $F(A,B)$ has nonempty interior [2]. This theorem also holds in higher dimensions, see [10, Chapter II] and the references therein and [29, Proposition 2.9 (i)].
@@
A topological analogue of Steinhaus’ theorem known as Piccard’s theorem asserts that the sum of two non-meager Baire sets (see Definition 1.2) in a topological vector space contains a nonempty open set [24, 16]. Again, addition can be replaced by a two variable function satisfying minimal differentiability assumptions; see [29, Proposition 2.9 (iii)] and [10, Chapter II]. There is also an analogous statement for topological groups, which states that $A*B$ (and $A*B^{-1}$) contains a nonempty open set provided that $A$ and $B$ are non-meager Baire sets [25].
While the aforementioned results consider operations on pairs of points, the present article focuses on general patterns and universality. A set or pattern $P$ is universal in the collection of sets $\mathcal{X}$ if for all $A\in\mathcal{X}$, we can find an affine copy of $P$ inside of $A$.
Universality can be considered for different classes of sets. We call a set measure universal if it is universal in the collection of Lebesgue measurable sets. Finite sets are known to be measure universal by Steinhaus’ theorem. Further, we say that $P$ is full measure universal if it is universal in the collection of measurable sets $A$ with the property that the compliment of $A$ has measure zero. Measure universality implies full measure universality. However, while it is easy to see that countably infinite sets are full measure universal (see Example A.6), it is an open question whether there exists such a set that is measure universal.
The classic Erdős similarity conjecture asserts that no infinite sets can be measure universal. While there are several examples of bounded countably infinite sets that are not measure universal, it is an open question whether there exists a countably infinite sequence that is measure universal. For a summary of results on recent variants of this conjecture of Erdős, see the survey of Jung, Lai, and Mooroogen [13].
Erdős’ similarity conjecture can also be studied in a topological setting. Here, Baire sets are the analogue of Lebesgue measurable sets, non-meager sets are the analogue of positive measure sets, and comeager sets are the analogue of full measure sets (see Definition 1.2 and Table 1.1).
Universality for the collection of comeager sets is completely resolved, which was recently considered by Gallagher, Lai and Weber [5]. Indeed, a set is universal among comeager sets if and only if it is a strong measure zero set [7] (also see [12]). Since Borel’s conjecture, which states that all strong measure zero sets are countable, is known to be independent of ZFC [28], it follows that the existence of an uncountable set that is universal among comeager sets is independent of ZFC.
In this article we provide the missing link for the topological variant of the classic Erdős similarity conjecture. We show that all bounded countable sets are universal in the collection of non-meager Baire sets. Other than results inferred from the obvious implication that lack of comeager universality implies lack of non-meager universality, no previous results on non-meager universality have appeared in the literature to the authors’ knowledge.
@@
To put this into further context, it is a straightforward consequence of the Baire category theorem that countably infinite sets are universal in the collection of comeager sets (see Example A.6). Our main Euclidean result (Theorem 2.1) extends this observation to the collection of non-meager Baire sets. Further, we prove a more quantitative result about an interval worth of scalings, and our results extend to topological vector spaces (Theorem 3.3).
To end, we pose the following direct topological analogue of the Erdős similarity conjecture:
Conjecture 1.1.
For each uncountable set $P$, there exists a non-meager Baire set $A$ that contains no affine image of $P$.
@@ 1.1. Background and definitions
Let $X$ denote a topological space. A subset of $X$ is meager (first category) if it is a countable union of nowhere dense sets. A set is non-meager (second category) if it is not meager, and it is comeager if its complement is meager. Finally, a set $A\subset X$ is said to be Baire if there is an open set $U$ and a meager set $M$ so that $A$ equals the symmetric difference $A=U\Delta M$.
An example of a set that is non-meager but not comeager is $[0,1]$; more generally, any set with both nonempty interior and exterior is a non-meager set which is not comeager.
@@
Equivalently, one may define Baire sets as sets $A$ for which there exists an open set $U$ such that $U\Delta A$ is meager. If $A$ satisfies this condition, then $A$ satisfies Definition 1.2 with $P=U\Delta A$. Conversely, if $U$ satisfies Definition 1.2, then $A\Delta U=M$ is meager. See Theorem A.1 for other useful equivalent definitions.
Measure notions Topological notions
sets of full measure comeager sets
sets of positive measure non-meager sets
measurable sets Baire sets
measure zero sets meager sets
Morally, one may think of meager sets as the topological analogues of measure zero sets (in the sense that they are considered negligible and are stable under countable unions), and of Baire sets as the topological analogues of measurable sets (in the sense that they are “almost” equal to open sets up to negligible error). A word of caution: even meager sets can have full measure. A succinct account of basic facts and examples is given in Appendix A; for a more thorough introduction, see [14, Chapter I.8].
@@
The definition of $\Delta_{P}(A)$ given above immediately extends to vector spaces:
@@
Let $V$ be a real vector space. Given sets $A,P\subset V$, define
$\Delta_{P}(A)=\{t\in\mathbb{R}:\exists z\in V\ tP+z\subset A\}.$
In the notation above, a consequence of Piccard’s Theorem is that $\Delta_{P}(A)$ has nonempty interior whenever $A$ is a non-meager Baire set and $P$ is a two-point set. In the following sections, we explore generalizations of this result to more complex point configurations and more abstract spaces. To make sense of the set $\Delta_{P}(A)$, we need both translations and scalings, and for this reason we work in topological vector spaces.
@@ 1.2. Acknowledgment
K.T. is supported in part by the Simons Foundation Grant GR137264. The authors thank Angel Cruz, Yeonwook Jung, Yuveshen Mooroogen, and the anonymous referee for interesting discussions and suggestions related to this article.
@@ 2. Results in Euclidean space
Let $A\subset\mathbb{R}^{d}$ be a non-meager Baire set, and let $P=\{v^{i}\}_{i\in\mathbb{N}}$ be a bounded sequence. Then, $\Delta_{P}(A)$ has nonempty interior.
@@
Since $A$ is Baire, there is an open set $U$ such that $A\Delta U$ is meager (since $A=U\Delta M$ implies $A\Delta U\subset M$). Since $A$ is non-meager, $U$ must be nonempty. Let $B$ be any open ball contained in $U$. Since
@@
is meager. If $A\cap B_{t}$ were meager, it would follow that $(I\cap J_{t})$ is meager. This cannot happen if $t$ is chosen so that $I\cap J_{t}$ is a non-degenerate interval. Denote the set of such $t$ by $K=\{t:\left(I\cap J_{t}\right)^{\circ}\neq\emptyset\},$ and observe that $K$ is a non-degenerate interval. Therefore, for any $t\in K$, we have $A\cap B_{t}\neq\varnothing$. ∎
This proof idea generalizes to arbitrary countable sets without too much difficulty. The key geometric fact we need to use is that the intersection of infinitely many balls still has nonempty interior, provided that each ball is a small translate of a single original ball.
@@
Define $A_{i,t}=A-tv^{i}$. We have $\{tv^{i}+z:i\in\mathbb{N}\}\subset A$ if and only if $z\in\bigcap_{i=1}^{\infty}A_{i,t}$. We must show this intersection is nonempty for an interval worth of $t$. Apply Lemma 2.2 to obtain an open ball $B$ with $B\setminus A$ meager, and let $B_{i,t}=B-tv^{i}$. It follows that $B_{i,t}\setminus A_{i,t}$ is meager for every $i,t$. Finally, let
$\widetilde{B}_{t}=\bigcap_{i=1}^{\infty}B_{i,t}.$
The set $\widetilde{B}_{t}\setminus A_{i,t}$ is meager, since $\widetilde{B}_{t}\subset B_{n,i}$. Therefore,
$\widetilde{B}_{t}\setminus\left(\bigcap_{i=1}^{\infty}A_{i,t}\right)=\bigcup_{i=1}^{\infty}(\widetilde{B}_{t}\setminus A_{i,t})$
is meager. If $\bigcap_{i=1}^{\infty}A_{i,t}$ were meager, then $\widetilde{B}_{t}$ would be meager. However, if $t$ is sufficiently small (relative to the radius of $B$ and $\sup_{i}|v^{i}|$, which is finite by assumption), then each set $B_{i,t}$ is a small translate of the original ball $B$, and hence the intersection $\widetilde{B}_{t}$ has nonempty interior. Therefore, the theorem holds when $J$ is a sufficiently small neighborhood of 00. ∎
@@ 3. Results in topological vector spaces
We establish the claim for $A-B$. Let $U_{1}$, $U_{2}$ denote open sets, and let $M_{1}$, $M_{2}$ denote meager sets so that $A=U_{1}\Delta M_{1}$ and $B=U_{2}\Delta M_{2}$. Let us assume that nonempty open sets are not meager, for otherwise $X$ would also be meager by the Banach category theorem (see Theorem 3.4) as a union of translates of such a set.
@@
(4) $A\cap\left(t+B\right)\supset(U_{1}\cap(t+U_{2}))\backslash(M_{1}\cup(M_{2}+t)),$
and since $(U_{1}\cap(t+U_{2}))$ is a nonempty open set, while $(M_{1}\cup(M_{2}+t))$ is meager, it follows that the right-hand-side of (4) is nonempty.
Moving on, recall that in Theorem 2.1, our hypotheses included that the configuration set under consideration was bounded. Since an arbitrary topological vector space need not be metric, we need a definition of boundedness in this context.
@@
Let $V$ be a topological vector space, and let $P\subset V$. We say $P$ is bounded if, for any open set $U\subset V$, there exist $t>0,z\in V$ such that $tP+z\subset U$.
@@
Let $V$ be a topological vector space, and let $P=\{v^{n}\}_{n\in\mathbb{N}}\subset V$ be a bounded sequence. If $A\subset V$ is a non-meager Baire set, then $\Delta_{P}(A)$ has nonempty interior.
@@
See [23, Chapter 16] for an exposition of this result. As an immediate consequence, we have:
@@
Next, we need a topological vector space analogue of the open balls of Euclidean space. In Definition 3.2, if $P$ is a bounded set and we scale $P$ by $t_{0}$ to be in a given neighborhood of 0, then there is no guarantee that we could then scale by a smaller $t<t_{0}$ and still guarantee this containment. Fortunately, in any topological vector space, there is a topological base consisting of sets which are closed under scaling down. This is captured in the following lemma.
@@
Let $V$ be a topological vector space, let $U$ be an open set, and let $P$ be a bounded set. For each $x\in P$, define $U_{x,t}=U-tx$. Finally, define
$\widetilde{U}_{t}=\bigcap_{x\in P}U_{x,t}.$
@@
is open. Since Cartesian products of open sets are a basis for the product topology, there exist open sets $U_{1},U_{2}\subset V$ such that $(\textbf{0},\textbf{0})\in U_{1}\times U_{2}$ and $U_{1}+U_{2}\subset U$. The set $U^{\prime}=U_{1}\cap U_{2}\cap U$ is therefore an open neighborhood of 0, contained in $U$, with the property that $U^{\prime}+U^{\prime}\subset U$. By Lemma 3.7, we may take a subset $B\subset U^{\prime}$ with these same properties, which is also balanced. Since $P$ is bounded, there exists $t_{0}$ such that $t_{0}P\subset B$. Since $B$ is balanced, we therefore have $tP\subset B$ for any $t\in J:=(0,t_{0})$. We claim that for any such $t$, the set $\widetilde{U}_{t}$ contains $B$ as a subset. Indeed, for any $z\in B,t\in J$, and $x\in P$, we have
@@
Define $A_{i,t}=A-tv^{i}$. We have $\{tv^{i}+z:i\in\mathbb{N}\}\subset A$ if and only if $z\in\bigcap_{i=1}^{\infty}A_{i,t}$. We must show that this intersection is nonempty for an interval worth of $t$. Let $U$ be an open set, and let $M$ be a meager set such that $A=U\Delta M$; let $U_{i,t}=U-tv^{i}$ and $M_{i,t}=M-tv^{i}$. Since any translation of a nowhere dense set is still nowhere dense, the set $M_{i,t}$ is meager for every $i,t$. Define
$\widetilde{U}_{t}=\bigcap_{i=1}^{\infty}U_{i,t}\,\,\,\text{ and }\,\,\,\widetilde{M}_{t}=\bigcup_{i=1}^{\infty}M_{i,t}.$
@@
$\bigcap_{i=1}^{\infty}A_{i,t}\supset\widetilde{U}_{t}\backslash\widetilde{M}_{t}.$
The proof is completed by arguing that $\widetilde{U}_{t}$ is non-meager; indeed, since $\widetilde{M}_{t}$ is meager, then it would follow that $\bigcap_{i=1}^{\infty}A_{i,t}$ is nonempty. By Lemma 3.8, there is a non-degenerate interval $J\subset\mathbb{R}$ such that for any $t\in J$, the set $\widetilde{U}_{t}$ has nonempty interior. By Lemma 3.5, since $V$ contains a non-meager set $A$, a set with nonempty interior must be non-meager.